# To Label or Not to Label: Hybrid Active Learning for Neural Machine Translation

### v1

@@ 4.3 Baselines
AL strategies for NMT primarily include a sentence-level selection of examples or phrase-level selection for training the NMT model. Sentence-level strategies focus on selecting sentences that are most useful for training while phrase-level strategies choose individual phrases that are the most informative. In some scenarios, phrase-level strategies can be more cost-effective as they provide granular control to avoid the inclusion of otherwise informative sentences which have phrases that the NMT model can already translate well. However, their implementation for NMT is generally more complex than sentence-level selection, involving the creation of synthetic parallel data that incorporates phrasal translations followed by a data mixing procedure that is utilized for mixed fine-tuning Hu and Neubig (2021). Thus, we compare our approach to three strong sentence-level selection baselines, including Random selection, Normalized Sequence Probability Ueffing and Ney (2007) and In-Domain Diversity Sampling Tsvigun et al. (2022a). The extension of the hybrid sampling strategy to phrase-level selection is left for future work.
While BADGE and ALPS are promising AL methods for classification tasks, applying them to NMT requires significant modification to the acquisition functions. Thus, we do not include them as baselines for AL in NMT.
4.3.1 Random Sampling
@@
Random sampling involves selecting a subset of sentences randomly from the unannotated data for labeling. Despite its simplicity, it has shown strong performance in active learning for various NLP tasks Miura et al. (2016); Tsvigun et al. (2022a) as it reflects the distribution of the complete unlabeled set without bias. Moreover, it is an efficient strategy as no queries are needed for example selection in contrast to the other AL strategies.
4.3.2 Normalized Sequence Probability (NSP)
@@
This is an uncertainty-based query strategy that helps to quantify the uncertainty of a model’s predictions on a particular sequence. It was introduced by Ueffing and Ney (2007) and is the basis for several other AL techniques Haffari et al. (2009). NSP for a sentence $\mathbf{v}$ is defined as:
$NSP(\mathbf{v})=1-\bar{p}_{\hat{\mathbf{w}}}(\mathbf{y}\mid\mathbf{v})$ (7)
where $\bar{p}_{\hat{\mathbf{w}}}(\mathbf{y}\mid\mathbf{v})$ is the geometric mean of the token probabilities predicted by model.
4.3.3 In-Domain Diversity Sampling (IDDS)
@@
This strategy Tsvigun et al. (2022a) attempts to select sentences that are different from the previously annotated ones in an attempt to increase the diversity of the final set. Simultaneously, it prevents the selection of noisier instances (or outliers) that are significantly different semantically from the documents in the in-domain dataset. This mitigates the prevalent issue of selecting noisier instances found in other uncertainty-based AL strategies. The acquisition function of IDDS is given by:
@@ 5 Results
Figure 2 shows the BLEU scores of BART-base model fine-tuned with different AL strategies across three datasets. We do three runs with different random seeds for each selection strategy and dataset and report the mean SacreBLEU. The shaded region shows the standard deviation across runs for a particular strategy and dataset. We observe that HUDS performs better than random sampling, NSP, and IDDS across all iterations. For the medicine dataset, HUDS demonstrates 34.2 SacreBLEU at the end of the final iteration compared to 33.6 for NSP, 33.3 for IDDS, and 33.1 for Random. We also notice that the relative ranking of SacreBLEU for different strategies varies significantly across datasets. We hypothesize that this indicates a difference in the distribution and, more specifically, the uncertainty and diversity of instances that we uncover next. Interestingly, IDDS does not perform well on the IT dataset, with a 13.8% lower SacreBLEU compared to random sampling (25.4 vs 29.5). We conjecture that this is due to the selection of diverse instances in each iteration that are different from the already annotated ones but are uninformative for the model.
Figure 3: Plot of uncertainty vs diversity scores for Medicine (Row 1), Law (Row 2), and IT (Row 3) datasets. Each plot contains 100 markers representing sentences that were selected during the AL procedure over 10 iterations, with 10 sentences selected in each iteration. The color represents sampling iteration in AL, with the darkest color representing the first iteration and the lightest color representing the tenth iteration. Columns correspond to the sampling strategy. Across all three datasets, hybrid sampling tends to select sentences with high uncertainty and diversity, while uncertainty and diversity sampling prefer sentences with high uncertainty scores and high diversity scores respectively.
@@ 5.1 Analyzing HUDS
To understand the sentence selection preference in HUDS during AL iterations, it is important to find out how it compares to regular uncertainty and diversity sampling. We plot uncertainty vs. diversity scores (Fig. 3) and analyze the unlabeled instances selected in ten iterations for annotation by each AL strategy (uncertainty, diversity, HUDS) across three datasets. In the earlier iterations (represented by dark-colored shapes), HUDS selects instances with high values of uncertainty and diversity for Medicine and IT datasets. For the Medicine and Law datasets, a few diverse instances have moderate to low uncertainty (Row 1 and 3, Column 2 of Fig. 3); hence HUDS selects those as well. In contrast, uncertainty sampling always selects uncertain instances regardless of their diversity, and diversity sampling prefers heterogeneous instances only. The difference in the distribution of uncertainty and diversity across datasets also helps explain the difference in the relative ranking of strategies in terms of SacreBLEU in Fig. 2.
@@ 5.2 Composition of annotated sentences
To find out how the composition of annotated instances affects the generalization performance of the NMT model, it is useful to analyze the vocabulary overlap between the selected sentences and the test set. Thus, we conduct an experiment to analyze the uni-gram overlap between the instances selected by different methods and the validation set of IT corpus, following the protocol presented by Hu and Neubig (2021). We find that the sentences selected by HUDS have the highest overall uni-gram overlap with the validation set, followed by random, NSP, and IDDS (Table 2). This indicates that HUDS selects varied instances with higher uni-gram coverage, potentially leading to better performance. We also compare the overlap of English and German individually, which shows a similar trend, except that NSP has a higher English unigram overlap than the randomly annotated sentences (46.7 vs 46.6).
@@
Table 2: Percentage overlap between the uni-grams of sentences selected by different AL strategies and the validation set of IT corpus. We report the individual overlap for En and De and the En-De combined overlap. HUDS demonstrates the highest En-De overlap, followed by NSP, Random and IDDS.
5.3 Ablation study
An important hyperparameter in the HUDS algorithm is the parameter $\lambda$ that balances uncertainty and diversity score for a sentence. We now study the effect of varying the value of $\lambda$ in HUDS on the SacreBLEU score for the multi-domains dataset. We run a separate AL procedure for each value of $\lambda$ on different datasets and report SacreBLEU score at the end of the first iteration with 1000 sentences annotated and used for training in that iteration. The results in Table 3 demonstrate a higher SacreBLEU score for $\lambda=0.5$, indicating that a balance between hybrid uncertainty and diversity scores in HUDS leads to better generalization. Additionally, the parameter $\lambda$ provides a powerful tuning method for uncertainty or diversity sampling using just the validation set.
@@ 6 Related Work
6.1 Active Learning in Machine Translation
Active Learning has been widely applied to machine translation (MT) to improve the quality and efficiency of translation systems, especially for low-resource languages or domains Zhang et al. (2018); Hu and Neubig (2021); Gupta et al. (2021); Zhou and Waibel (2021); Zhao et al. (2020). Several aspects of AL have been explored for MT, including different sampling techniques that consider uncertainty or diversity Ambati et al. (2011), different workflows including batch mode vs. online mode Ananthakrishnan et al. (2010); Lam et al. (2019), leveraging pre-trained or auxiliary models Domhan and Hieber (2017); Wang et al. (2020) and exploiting domain-specific and synthetic data Zhang et al. (2019); Hoang et al. (2018).
6.2 Uncertainty Sampling Techniques
Some of the early works on AL for MT use uncertainty-based sampling criteria, such as entropy or posterior probability, to select the sentences that the model is most uncertain about for annotation Haffari et al. (2009); Bloodgood and Vijay-Shanker (2009); Zhao et al. (2020). Using these sentences will enable the model to generalize better to unseen data. For NMT models, uncertainty is usually measured at the sentence or the beam level, based on the output of an encoder-decoder model or a transformer model Zhou et al. (2020). However, pure uncertainty sampling techniques suffer from redundancy issues as they tend to select uncertain similar sentences or prefer certain words and structures.
6.3 Diversity Sampling Techniques
The main idea of diversity sampling is to select the sentences that are most diverse and representative of the unlabeled data based on some diversity measures such as density, representativeness, or coverage Bloodgood and Callison-Burch (2010); Gangadharaiah et al. (2009); Hu and Neubig (2021); Pendas et al. (2023). Diversity is usually measured at the word or phrase level based on the frequency or similarity of words or phrases. However, pure diversity sampling methods lead to the selection of noisier or irrelevant instances that are too difficult or too easy for the model.
@@ 7 Limitations
We list the potential limitations of our work below:
  1. 1.
    HUDS involves computation of embeddings, clustering, and calculation of hybrid sampling scores for each unlabeled sentence, which could lead to higher latency compared to other AL methods. For large datasets, the additional computations may affect the interactivity of the AL procedure. Precomputation of embeddings and intermediate caching of scores can be leveraged to mitigate this.
  2. 2.
    The proposed hybrid sampling strategy HUDS focuses on sentence-level active learning for NMT. Phrase-level active learning could potentially be more cost-effective by selecting only informative phrases within sentences. Extending HUDS to phrase-level selection is left for future work.
8 Conclusion
In this work, we proposed HUDS, a novel hybrid active learning strategy that leverages both uncertainty and diversity for neural machine translation. Experimental results on three domain-specific machine translation datasets showed that HUDS outperforms other AL methods, demonstrating the effectiveness of hybrid sampling for active learning in NMT. Future work includes (1) incorporating phrasal selection in HUDS, (2) evaluating HUDS on low-resource languages where it can potentially enhance the performance of large language models using minimal data, and (3) making HUDS more efficient through precomputation of embeddings and intermediate caching of scores.
@@ A.2 Datasets
We use the WMT14 En-De dataset available under the CC-BY-SA license and the German-English multi-domains parallel data.
@@ Appendix C Resources
The experiments are run on a single 48GB NVIDIA A6000 GPU on a privately hosted server. Around 1350 GPU hours were used for the entirety of the project which includes the early experiments and the final results.

### v2

@@ 4.3 Baselines
AL strategies for NMT primarily include a sentence-level selection of examples or phrase-level selection for training the NMT model. Sentence-level strategies focus on selecting sentences that are most useful for training while phrase-level strategies choose individual phrases that are the most informative. In some scenarios, phrase-level strategies can be more cost-effective as they provide granular control to avoid the inclusion of otherwise informative sentences which have phrases that the NMT model can already translate well. However, their implementation for NMT is generally more complex than sentence-level selection, involving the creation of synthetic parallel data that incorporates phrasal translations followed by a data mixing procedure that is utilized for mixed fine-tuning Hu and Neubig (2021a). Thus, we compare our approach to three sentence-level selection baselines, including Random selection, Normalized Sequence Probability (NSP) Ueffing and Ney (2007), and In-Domain Diversity Sampling (IDDS) Tsvigun et al. (2022a). NSP and IDDS are strong baselines, while random selection is competitive in AL given a sufficiently large annotation budget Hu and Neubig (2021b).
While BADGE and ALPS are promising AL methods for classification tasks, applying them to NMT requires significant modification to the acquisition functions. Thus, we do not consider them as baselines for AL in NMT. Additionally, we do not include subset selection approaches as baselines, e.g., Kothawade et al. (2022), as their goal is to select a subset from a fixed set of (often fully labeled) examples to allow training in resource-constrained environments. In contrast, active learning aims to iteratively select a subset of examples for labeling from an unlabeled pool, with the aim of minimizing the labeling cost.
@@
Random Sampling. Random sampling involves selecting a subset of sentences randomly from the unannotated data for labeling. Despite its simplicity, it has shown strong performance in active learning for various NLP tasks Miura et al. (2016); Tsvigun et al. (2022a) as it reflects the distribution of the complete unlabeled set without bias. Moreover, it is an efficient strategy as no queries are needed for example selection in contrast to the other AL strategies.
@@
Normalized Sequence Probability (NSP). This is an uncertainty-based query strategy that helps to quantify the uncertainty of a model’s predictions on a particular sequence. It was introduced by Ueffing and Ney (2007) and is a strong baseline that is the basis for several other AL techniques Haffari et al. (2009); Tsvigun et al. (2022b). NSP for a sentence $\mathbf{v}$ is defined as $1-\bar{p}_{\hat{\mathbf{w}}}(\mathbf{y}\mid\mathbf{v})$, where $\bar{p}_{\hat{\mathbf{w}}}(\mathbf{y}\mid\mathbf{v})$ is the geometric mean of the token probabilities predicted by model.
@@
In-Domain Diversity Sampling (IDDS). This strategy Tsvigun et al. (2022a) attempts to select sentences that are different from the previously annotated ones in an attempt to increase the diversity of the final set. Simultaneously, it prevents the selection of noisier instances (or outliers) that are significantly different semantically from the documents in the in-domain dataset. This mitigates the prevalent issue of selecting noisier instances found in other uncertainty-based AL strategies. The acquisition function of IDDS is given by:
@@
Figure 3: Plot of uncertainty vs diversity scores for Medicine (Row 1), Law (Row 2), and IT (Row 3) datasets. Each plot contains 100 markers representing sentences that were selected during the AL procedure over 10 iterations, with 10 sentences selected in each iteration. The color represents sampling iteration in AL, with the darkest color representing the first iteration and the lightest color representing the tenth iteration. Columns correspond to the sampling strategy. Across all three datasets, hybrid sampling tends to select sentences with high uncertainty and diversity.
@@ 5 Results
Table 2 shows the BLEU scores of BART-base model fine-tuned with different AL strategies across three datasets. We do three runs with different random seeds for each selection strategy and dataset and report the mean SacreBLEU. We observe that HUDS performs better than random sampling, NSP, and IDDS across all iterations. For the medicine dataset, HUDS demonstrates 34.58 SacreBLEU at the end of the final iteration compared to 33.68 for NSP, 33.05 for IDDS, and 32.70 for Random. The differences in the SacreBLEU scores are statistically significant by Wilcoxon’s signed-rank test (p < 0.05). We also notice that the relative ranking of SacreBLEU for different strategies varies significantly across datasets. We hypothesize that this indicates a difference in the distribution and, more specifically, the uncertainty and diversity of instances that we uncover next. Interestingly, IDDS does not perform well on the IT dataset, with a 13.8% lower SacreBLEU compared to random sampling (25.4 vs 29.5). We conjecture that this is due to the selection of diverse instances in each iteration that are different from the already annotated ones but are uninformative for the model.
@@ 5.1 Analyzing HUDS
To understand the sentence selection preference in HUDS during AL iterations, it is important to find out how it compares to regular uncertainty and diversity sampling. We plot uncertainty vs. diversity scores (Fig. 3) and analyze the unlabeled instances selected in ten iterations for annotation by each AL strategy (uncertainty, diversity, HUDS) across three datasets. In the earlier iterations (represented by dark-colored shapes), HUDS selects instances with high values of uncertainty and diversity for Medicine and IT datasets. For the Medicine and Law datasets, a few diverse instances have moderate to low uncertainty (Row 1 and 3, Column 2 of Fig. 3); hence HUDS selects those as well. In contrast, uncertainty sampling always selects uncertain instances regardless of their diversity, and diversity sampling prefers heterogeneous instances only. The difference in the distribution of uncertainty and diversity across datasets also helps explain the difference in the relative ranking of strategies in terms of SacreBLEU in Table 2.
5.2 Evaluation on En-Fr corpus
Random IDDS NSP HUDS
34.8 33.5 35.1 35.5
Table 3: Comparison of the mean SacreBLEU after the final AL iteration on the IWSLT 2014 En-Fr validation set.
To confirm the generalization of HUDS to other languages, we conduct an experiment on the English-French (En-Fr) language pair. We utilize the WMT14 (En-Fr) parallel data as the out-of-domain labeled dataset and the IWSLT 2014 Ted Talks (En-Fr) as the in-domain dataset, following the protocol presented in Wang et al. (2017). We use the BART-base model for AL. HUDS demonstrates the highest SacreBLEU score on the IWSLT 2014 validation set compared to other strategies, confirming its generalization to En-Fr (Table 3).
5.3 Changing annotation size
To determine the impact of a larger annotation size on the translation quality, we conduct an experiment on the IT dataset with 5000 examples labeled in each iteration (instead of 1000 in the earlier experiments). Ten iterations are performed, which brings the final quantity of selected sentences to 50,000. The results in Table 4 show that HUDS consistently shows a higher SacreBLEU score than other methods with a larger annotation size.
Method Iteration
2 4 6 8 10
Random 29.7 30.7 31.5 31.9 32.5
IDDS 25.6 27.1 27.9 28.7 29.1
NSP 30.5 31.2 31.9 32.0 32.3
HUDS 30.8 32.3 32.9 33.4 34.0
Table 4: Mean SacreBLEU score on the IT dataset. 5000 examples are selected for labeling in each iteration.
@@ 5.4 Composition of annotated sentences
To find out how the composition of annotated instances affects the generalization performance of the NMT model, it is useful to analyze the n-gram overlap between the selected sentences and the test set. Prior work Hu and Neubig (2021a) has shown that the n-gram overlap correlates highly with test BLEU scores, making it a reliable measure of the usefulness of selected data for domain adaptation. Thus, we conduct an experiment to analyze the unigram overlap between the instances selected by different methods and the validation set of IT corpus, following the protocol presented in Hu and Neubig (2021a). We find that the sentences selected by HUDS have the highest overall unigram overlap with the validation set, followed by random, NSP, and IDDS (Table 5). This indicates that HUDS selects varied instances with higher unigram coverage, potentially leading to better performance. We also compare the overlap of English and German individually, which shows a similar trend, except that NSP has a higher English unigram overlap than the randomly annotated sentences (46.7 vs 46.6).
@@
Table 5: Percentage overlap between the unigrams of sentences selected by different AL strategies and the validation set of IT corpus. We report the individual overlap for En and De and the En-De combined overlap.
5.5 COMET scores
We additionally assess the translation quality of HUDS using COMET Rei et al. (2020), a state-of-the-art neural evaluation metric that demonstrates high correlation with human judgment. We use the Unbabel/wmt20-comet-da model [^2] https://github.com/Unbabel/COMET for scoring the translations on the IT domain dataset. The results in Table 6 show that HUDS outperforms other active learning strategies.
Random IDDS NSP HUDS
42.5 38.0 42.3 42.6
Table 6: Comparison of the COMET scores after the final AL iteration for different AL strategies on the medical domain dataset. Unbabel/wmt20-comet-da model is used for evaluation. Active learning is done over ten iterations, with 1000 sentences selected in each iteration.
6 Ablation study
An important hyperparameter in the HUDS algorithm is the parameter $\lambda$ that balances uncertainty and diversity score for a sentence. We now study the effect of varying the value of $\lambda$ in HUDS on the SacreBLEU score for the multi-domains dataset. We run a separate AL procedure for each value of $\lambda$ on different datasets and report SacreBLEU score at the end of the first iteration with 1000 sentences annotated and used for training in that iteration. The ablation is done on a single iteration to better isolate the impact of individual components and avoid any potential confounding effects. The results in Table 7 demonstrate a higher SacreBLEU score for $\lambda=0.5$, indicating that a balance between hybrid uncertainty and diversity scores in HUDS leads to better generalization. Additionally, the parameter $\lambda$ provides a powerful tuning method for uncertainty or diversity sampling using just the validation set.
Dataset $\lambda$
0 0.1 0.3 0.5 0.7 0.9 1.0
Medicine 29.6 29.6 29.7 29.9 29.7 29.7 29.5
Law 32.8 32.8 32.7 33.0 32.8 32.9 31.2
IT 28.7 28.8 28.6 29.0 28.7 28.6 28.5
@@ 7 Related Work
Active Learning in Machine Translation. Active Learning has been widely applied to machine translation (MT) to improve the quality and efficiency of translation systems, especially for low-resource languages or domains Zhang et al. (2018); Hu and Neubig (2021a); Gupta et al. (2021); Zhou and Waibel (2021); Zhao et al. (2020). Several aspects of AL have been explored for MT, including different sampling techniques that consider uncertainty or diversity Ambati et al. (2011), different workflows including batch mode vs. online mode Ananthakrishnan et al. (2010); Lam et al. (2019), leveraging pre-trained or auxiliary models Domhan and Hieber (2017); Wang et al. (2020) and exploiting domain-specific and synthetic data Zhang et al. (2019); Hoang et al. (2018). To our knowledge, no hybrid AL strategy for efficiently acquiring domain-specific data in NMT has been proposed prior to our work.
Uncertainty Sampling Techniques. Some of the early works on AL for MT use uncertainty-based sampling criteria, such as entropy or posterior probability, to select the sentences that the model is most uncertain about for annotation Haffari et al. (2009); Bloodgood and Vijay-Shanker (2009); Zhao et al. (2020). Using these sentences will enable the model to generalize better to unseen data. For NMT models, uncertainty is usually measured at the sentence or the beam level, based on the output of an encoder-decoder model or a transformer model Zhou et al. (2020). However, pure uncertainty sampling techniques suffer from redundancy issues as they tend to select uncertain similar sentences or prefer certain words and structures.
Diversity Sampling Techniques. The main idea of diversity sampling is to select the sentences that are most diverse and representative of the unlabeled data based on some diversity measures such as density, representativeness, or coverage Bloodgood and Callison-Burch (2010); Gangadharaiah et al. (2009); Hu and Neubig (2021a); Pendas et al. (2023). Diversity is usually measured at the word or phrase level based on the frequency or similarity of words or phrases. However, pure diversity sampling methods lead to the selection of noisier or irrelevant instances that are too difficult or too easy for the model.
8 Conclusion
In this work, we proposed HUDS, a novel hybrid active learning strategy that leverages both uncertainty and diversity for neural machine translation. HUDS outperforms other AL methods for NMT, demonstrating the effectiveness of hybrid sampling over other baselines for the same amount of human labeling effort. Future work includes (1) incorporating phrasal selection in HUDS and (2) making HUDS more efficient through precomputation of embeddings and intermediate caching of scores.
@@ 9 Limitations
We list the potential limitations of our work below: (1) HUDS involves computation of embeddings, clustering, and calculation of hybrid sampling scores for each unlabeled sentence, which could lead to higher latency compared to other AL methods. For large datasets, the additional computations may affect the interactivity of the AL procedure. Precomputation of embeddings and intermediate caching of scores can be leveraged to mitigate this. HUDS can also be augmented with automatic quality estimation methods for NMT Blain et al. (2023) to increase efficiency. (2) The proposed hybrid sampling strategy HUDS focuses on sentence-level active learning for NMT. Phrase-level active learning could potentially be more cost-effective by selecting only informative phrases within sentences. Extending HUDS to phrase-level selection is left for future work.
@@ A.2 Datasets
We use the WMT14 En-De dataset available under the CC-BY-SA license and the German-English multi-domains parallel data available under the Creative Commons CC0 license.
@@ Appendix C Resources
The experiments are run on a single 48GB NVIDIA A6000 GPU on the cloud. Around 1680 GPU hours were used for the entirety of the project which includes the early experiments and the final results.